If the perimeter of a rectangle is 42 cm and the area is 90 square cm, then what are the dimensions of the rectangle?

Answer by ewatrrr(24785)

 
Hi,

P = 2L +2w = 42cm
L + w = 21 L+=+21-w
A = Lw
(21-w)w = 90cm^2
w^2 - 21w + 90 = 0
(w-15)(w-6) = 0
Dimensions: 15cm by 6cm